The sum of three consecutive even numbers is 150. Find the numbers.
What does it mean consecutive even numbers? How do I do this?
Thank you

Answer by rapaljer(4551)

Consecutive even numbers means like 2,4, and 6.
Or 18, 20, 22.
Or 74, 76, 78.

In general the second number is always 2 more than the first one, and the third one is always 4 more than the first one.

So, let x= first even number
x+2= second even number
x+4= third even number.

If three consecutive even numbers is 150, then
x+ x+2+x+4= 150
3x+6=150
3x=144
x= 48
x+2=50
x+4=52

Check: The sum of the numbers 48+50+52=150. It checks!

For additional explanation on Word Problems, in particular "Consecutive Integer" problems, please see my own website. Do a "Bing" or "Google" search for my last name "Rapalje". Look for "Rapalje Homepage" near the top of these searches. Then near the top of my Homepage, look for "Basic, Intermediate and College Algebra". Choose "Basic Algebra", "Chapter 1", "Section 1.10 Word Problems Part I". See also the "MATH IN LIVING COLOR" pages that go with this section.

R^2

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus